OpenStudy (anonymous):

Find the f-1(x). State any restrictions on the domain of f-1(x). f(x)=x2-6x+1, x≥3

myininaya (myininaya):

you have y=x^2-6x+1 or you can write this as 0=x^2-6x+1-y Solve this for x.

myininaya (myininaya):

You could use the quadratic formula

OpenStudy (anonymous):

Thank you, this was a new one for me. I appreciate it.

myininaya (myininaya):

You will get two options for x use the fact that x is greater than or equal to 3 to choose the appropriate choice then when done replace x with f^(-1)(x) and replace y with x